# Sublinear Regret for a Class of Continuous-Time Linear–Quadratic Reinforcement Learning Problems

### v2

@@ Abstract
We study reinforcement learning (RL) for a class of continuous-time linear–quadratic (LQ) control problems for diffusions, where states are scalar-valued and running control rewards are absent but volatilities of the state processes depend on both state and control variables. We apply a model-free approach that relies neither on knowledge of model parameters nor on their estimations, and devise an actor–critic algorithm to learn the optimal policy parameter directly. Our main contributions include the introduction of an exploration schedule and a regret analysis of the proposed algorithm. We provide the convergence rate of the policy parameter to the optimal one, and prove that the algorithm achieves a regret bound of $O(N^{\frac{3}{4}})$ up to a logarithmic factor, where $N$ is the number of learning episodes. We conduct a simulation study to validate the theoretical results and demonstrate the effectiveness and reliability of the proposed algorithm. We also perform numerical comparisons between our method and those of the recent model-based stochastic LQ RL studies adapted to the state- and control-dependent volatility setting, demonstrating a better performance of the former in terms of regret bounds.
Keywords: C ontinuous-Time RL, Stochastic Linear–Quadratic Control, Model-Free Policy Gradient, Regret Bounds, Exploration Scheduling, Actor–Critic Algorithm
@@ 1 Introduction
Reinforcement learning (RL) has been developed to tackle complex control problems in largely unknown environments. Its successful applications range from strategic board games such as chess and Go (Silver et al.,, 2016, 2017) to robotic systems (Gu et al.,, 2017; Khan et al.,, 2020). However, RL has been predominantly studied for discrete-time, Markov decision processes (MDPs) with discrete state and control spaces, even though most real-life applications are inherently continuous-time with continuous state space and possibly continuous control space (e.g., autonomous driving, stock trading, and video game playing). More importantly, while one can turn a continuous-time problem into a discrete-time MDP upfront by time discretization, such an approach is very sensitive to time step size and performs poorly with small time steps Munos, (2006); Tallec et al., (2019); Park et al., (2021).
@@
While there were studies directly on continuous-time RL, these had been rare and far between (Baird,, 1994; Doya,, 2000; Vamvoudakis and Lewis,, 2010; Lee and Sutton,, 2021; Kim et al.,, 2021) up to just recent years, overall lacking a systematic and unified theory. Starting with Wang et al., (2020) that introduces an entropy-regularized relaxed control framework for continuous-time RL, a series of subsequent papers (Jia and Zhou, 2022a, ; Jia and Zhou, 2022b, ; Jia and Zhou,, 2023) develop theories on policy evaluation, policy gradient, and $q$-learning respectively within this framework. This strand of research is characterized by focusing on learning the optimal control policies directly without attempting to estimate or learn the model parameters, underlining a model-free (up to the unknown dynamics being governed by diffusion processes) and data-driven approach. The mathematical foundation of the entire theory is the martingale property of certain stochastic processes, the enforcement of which naturally leads to various temporal difference and actor–critic algorithms to train and learn $q$-functions, value functions, and optimal (stochastic) policies. Subsequently, there has been active follow-up research with various extensions and applications; see, e.g. Huang et al., (2022); Dai et al., (2023); Wang et al., (2023); Frikha et al., (2023); Wei and Yu, (2023).
@@
A crucial question in RL is the convergence and regret bounds of RL algorithms that provide theoretical guidance and guarantee their effectiveness and reliability. For LQ problems, such theoretical results exist, for example, for deterministic systems (Bradtke,, 1992; Fazel et al.,, 2018; Malik et al.,, 2019) as well as systems with identically and independently distributed noises (Abbasi-Yadkori and Szepesvári,, 2011; Abeille and Lazaric,, 2018; Cohen et al.,, 2018, 2019; Hambly et al.,, 2021; Wang et al.,, 2021; Cassel and Koren,, 2021; Yang et al.,, 2019; Zhou and Lu,, 2023; Chen et al.,, 2023; Simchowitz and Foster,, 2020), covering finite-horizon, infinite-horizon, and ergodic cases. These studies are nevertheless all for discrete-time models, with control not affecting the level of randomness in the state dynamics. Some of them, e.g. Abbasi-Yadkori and Szepesvári, (2011); Abeille and Lazaric, (2018); Hambly et al., (2021); Wang et al., (2021); Zhou and Lu, (2023), design their algorithms based on policy gradient. However, the gradient representations therein rely on estimations of the drift parameters; hence, the methods are essentially model-based. In addition, the semidefinite programming formulation in (Cohen et al.,, 2018, 2019) does not seem applicable to continuous-time systems.
@@
The algorithm proposed and analyzed in the present paper belongs to the class of actor–critic algorithms originally put forward by Konda and Tsitsiklis, (1999). Such algorithms for discrete-time systems have been studied in Wu et al., (2020); Xu et al., (2021); Cen et al., (2022), and in particular, for ergodic LQ problems in Yang et al., (2019); Chen et al., (2023) and for episodic linear MDPs in Cai et al., (2020); Zhong and Zhang, (2023). The “optimal” regret of these algorithms is mostly of the order $O(\sqrt{N})$, where $N$ is the number of episodes or timesteps. However, it is unclear whether they still work for the diffusion case where the volatility also depends on state and control. For general continuous-time diffusion environments, however, the aforementioned series of papers (Wang et al., (2020); Jia and Zhou, 2022a ; Jia and Zhou, 2022b ; Jia and Zhou, (2023)) and their follow-up study have not addressed the problems of convergence and regret bounds. These remain highly significant yet challenging open questions due to the model-free nature of the underlying approach and the stochastic approximation type of algorithms involved.
@@
Recently, there has been some progress on regret analysis for continuous-time stochastic LQ RL in (Basei et al.,, 2022; Szpruch et al.,, 2024) that achieve sublinear regrets of their respective algorithms. Both papers assume that the diffusion coefficients are constant independent of state and control, which is vital for their approaches to work. Again, in essence, these works are model-based because they apply either least-square or Bayesian methods to estimate model parameters and use the corresponding estimation errors to deduce the regret bounds. In particular, there is an intrinsic drawback in these model-based methods specific to LQ problems: optimal control policies are linear feedbacks of the state; hence, these methods may suffer from the unidentifiability problem. In addition, they require both the batch size and the number of timesteps to increase exponentially over episodes, adding substantial computational and memory costs.
@@
In this paper, we propose a policy gradient based actor–critic algorithm to solve a special class of continuous-time, finite-horizon stochastic LQ problems under the model-free, episodic RL setting, where the state processes are one dimensional and there is no running control award. Our main contributions are
@@
    We take a model-free approach to develop our algorithm, i.e., a policy gradient based (soft) actor–critic algorithm, and base our analysis on the stochastic approximation scheme. In particular, the policy gradient in this paper is a “model-free gradient” instead of a “model-based gradient” commonly taken in discrete-time RL. As a result, we do not need to estimate model primitives in the entire analysis, circumventing the issues discussed earlier arising from estimating/learning those model parameters.
@@
The remainder of the paper is structured as follows. Section 2 formulates the problem and provides some preliminary results necessary for the subsequent development. Section 3 describes and explains the steps leading to our RL algorithm. Section 4 presents the main theoretical results on convergence and a regret bound of the proposed algorithm. Section 5 reports the results of numerical experiments. Finally, Section 6 concludes. The appendices contain proofs of the main results and a detailed description of the numerical experiments.
@@ 2.1 Classical Stochastic LQ Control
$\mathrm{d}x^{u}(t)=(Ax^{u}(t)+B^{\top}u(t))\mathrm{d}t+\sum_{j=1}^{m}(C_{j}x^{u}(t)+D_{j}^{\top}u(t))\mathrm{d}W^{(j)}(t),$ (1)
@@
$\max_{u}\mathbb{E}\left[\int_{0}^{T}-\frac{1}{2}Qx^{u}(t)^{2}\mathrm{d}t-\frac{1}{2}Hx^{u}(T)^{2}\right],$ (2)
@@
$V_{CL}(t,x)=\max_{u}\mathbb{E}\left[\int_{t}^{T}-\frac{1}{2}Qx^{u}(s)^{2}\mathrm{d}s-\frac{1}{2}Hx^{u}(T)^{2}\Big{|}x^{u}(t)=x\right].$ (3)
@@
$V_{CL}(t,x)=-\frac{1}{2}\biggl{[}\frac{Q}{\Lambda}+(H-\frac{Q}{\Lambda})e^{\Lambda(t-T)}\biggl{]}x^{2},$ (4)
$u_{CL}(t,x)=-(\sum_{j=1}^{m}D_{j}D_{j}^{\top})^{-1}(B+\sum_{j=1}^{m}C_{j}D_{j})x,$ (5)
@@
We remark that if $x$ is of a higher dimension and/or there is a control running reward, then the optimal policy will depend on the solution of the differential Riccati equation and hence become time-dependent, for which our current method fails.
@@ 2.2 RL Theory for LQ
$\mathrm{d}x^{\pi}(t)=\widetilde{b}(x^{\pi}(t),\pi(\cdot,t))\mathrm{d}t+\sum_{j=1}^{m}\widetilde{\sigma}_{j}(x^{\pi}(t),\pi(\cdot,t))\mathrm{d}W^{(j)}(t),$ (6)
@@
$\widetilde{b}(x,\psi):=Ax+B^{\top}\int_{\mathbb{R}^{l}}u\psi(u)\mathrm{d}u,$ (7)
$\widetilde{\sigma}_{j}(x,\psi):=\sqrt{\int_{\mathbb{R}^{l}}(C_{j}x+D_{j}^{\top}u)^{2}\psi(u)\mathrm{d}u},\;\;(x,\psi)\in\mathbb{R}\times\mathcal{P}(\mathbb{R}^{l}).$ (8)
@@
$J(t,x;\pi)=\mathbb{E}\left[\int_{t}^{T}\left(-\frac{1}{2}Qx^{\pi}(s)^{2}+\gamma p^{\pi}(s)\right)\mathrm{d}s-\frac{1}{2}Hx^{\pi}(T)^{2}\Big{|}x^{\pi}(t)=x\right],$ (9)
@@
$V(t,x)=\max_{\pi}J(t,x;\pi).$ (10)
By Wang et al., (2020), the optimal value function and optimal randomized/stochastic (feedback) policy are determined by
$V(t,x)=-\frac{1}{2}k_{1}(t)x^{2}+k_{3}(t),$ (11)
$\pi(u\mid t,x)=\mathcal{N}\left(u\Big{|}-(\sum_{j=1}^{m}D_{j}D_{j}^{\top})^{-1}(B+\sum_{j=1}^{m}C_{j}D_{j})x,\frac{\gamma}{k_{1}(t)}(\sum_{j=1}^{m}D_{j}D_{j}^{\top})^{-1}\right),$ (12)
where $k_{1}>0$ and $k_{3}$ are certain functions of $t$ that can be determined completely by the model primitives, and $\mathcal{N}(\cdot|\mu,\Sigma)$ is the multivariate Gaussian density with mean $\mu$ and covariance $\Sigma$. These theoretical results cannot be used to compute the solution of the exploratory problem because the model parameters are unknown, yet they reveal the structure of the solution inherent to LQ RL (i.e. the optimal value function is quadratic in $x$ and optimal stochastic policy is Gaussian) that can be utilized to significantly reduce the complexity of function parameterization/approximation in learning.

### v4

@@ Abstract
We study reinforcement learning (RL) for a class of continuous-time linear–quadratic (LQ) control problems for diffusions, where states are scalar-valued and running control rewards are absent but volatilities of the state processes depend on both state and control variables. We apply a model-free approach that relies neither on knowledge of model parameters nor on their estimations, and devise an RL algorithm to learn the optimal policy parameter directly. Our main contributions include the introduction of an exploration schedule and a regret analysis of the proposed algorithm. We provide the convergence rate of the policy parameter to the optimal one, and prove that the algorithm achieves a regret bound of $O(N^{\frac{3}{4}})$ up to a logarithmic factor, where $N$ is the number of learning episodes. We conduct a simulation study to validate the theoretical results and demonstrate the effectiveness and reliability of the proposed algorithm. We also perform numerical comparisons between our method and those of the recent model-based stochastic LQ RL studies adapted to the state- and control-dependent volatility setting, demonstrating a better performance of the former in terms of regret bounds.
Keywords: Continuous-Time RL, Stochastic Linear–Quadratic Control, Model-Free Policy Gradient, Regret Bounds, Exploration Scheduling, RL Algorithm
@@ 1 Introduction
Reinforcement learning (RL) has been developed to tackle complex control problems in largely unknown environments. Its successful applications range from strategic board games such as chess and Go (Silver et al., , 2016, 2017) to robotic systems (Gu et al., , 2017; Khan et al., , 2020). However, RL has been predominantly studied for discrete-time, Markov decision processes (MDPs) with discrete state and control spaces, even though most real-life applications are inherently continuous-time with continuous state space and possibly continuous control space (e.g., autonomous driving, stock trading, and video game playing). More importantly, while one can turn a continuous-time problem into a discrete-time MDP upfront by time and state discretization, such an approach is very sensitive to time step size and performs poorly with small time steps Munos, (2006); Tallec et al., (2019); Park et al., (2021).
@@
While there were studies directly on continuous-time RL, these had been rare and far between (Baird, , 1994; Doya, , 2000; Vamvoudakis and Lewis, , 2010; Lee and Sutton, , 2021; Kim et al., , 2021) up to just recent years, overall lacking a systematic and unified theory. Starting with Wang et al., (2020) that introduces an entropy-regularized relaxed control framework for continuous-time RL, a series of subsequent papers (Jia and Zhou, 2022a, ; Jia and Zhou, 2022b, ; Jia and Zhou, , 2023) develop theories on policy evaluation, policy gradient, and $q$-learning respectively within this framework. This strand of research is characterized by focusing on learning the optimal control policies directly without attempting to estimate or learn the model parameters, underlining a model-free (up to the unknown dynamics being governed by diffusion processes) and data-driven approach. The mathematical foundation of the entire theory is the martingale property of certain stochastic processes, the enforcement of which naturally leads to various temporal difference and actor–critic type of RL algorithms to train and learn $q$-functions, value functions, and optimal (stochastic) policies. Subsequently, there has been active follow-up research with various extensions and applications; see, e.g. Huang et al., (2022); Dai et al., (2023); Wang et al., (2023); Frikha et al., (2023); Wei and Yu, (2025); Huang et al., (2024).
@@
A crucial question in RL is the convergence and regret bounds of RL algorithms that provide theoretical guidance and guarantee their effectiveness and reliability. For LQ problems, such theoretical results exist, for example, for deterministic systems (Bradtke, , 1992; Fazel et al., , 2018; Malik et al., , 2019) as well as systems with identically and independently distributed noises (Abbasi-Yadkori and Szepesvári, , 2011; Abeille and Lazaric, , 2018; Cohen et al., , 2018, 2019; Hambly et al., , 2021; Wang et al., , 2021; Yang et al., , 2019; Zhou and Lu, , 2023; Chen et al., , 2023; Simchowitz and Foster, , 2020; Cassel and Koren, , 2021), covering finite-horizon, infinite-horizon, and ergodic cases. These studies are nevertheless all for discrete-time models, with control not affecting the level of randomness in the state dynamics. Some of them, e.g. Abbasi-Yadkori and Szepesvári, (2011); Abeille and Lazaric, (2018); Hambly et al., (2021); Wang et al., (2021); Zhou and Lu, (2023), design their algorithms based on policy gradient. However, the gradient representations therein rely on estimations of the drift parameters; hence, the methods are essentially model-based. In addition, the semidefinite programming formulation in (Cohen et al., , 2018, 2019) does not seem applicable to continuous-time systems.
@@
The algorithm proposed and analyzed in the present paper belongs to the general class of actor–critic algorithms originally put forward by Konda and Tsitsiklis, (1999). Such algorithms for discrete-time systems have been studied in Wu et al., (2020); Xu et al., (2021); Cen et al., (2022), and in particular, for ergodic LQ problems in Yang et al., (2019); Chen et al., (2023) and for episodic linear MDPs in Cai et al., (2020); Zhong and Zhang, (2023). The “optimal” regret of these algorithms is mostly of the order $O(\sqrt{N})$, where $N$ is the number of episodes or timesteps. However, it is unclear whether they still work for the diffusion case where the volatility also depends on state and control. For general continuous-time diffusion environments, however, the aforementioned series of papers (Wang et al., (2020); Jia and Zhou, 2022a ; Jia and Zhou, 2022b ; Jia and Zhou, (2023)) and their follow-up study have not addressed the problems of convergence and regret bounds. Assuming oracle access to the policy gradient, Giegrich et al., (2024) establishes the exponential convergence of the policy gradient methods for the entropy regularized LQ problem. However, in our setting, the policy gradient needs to be estimated by samples using a model-free approach, leading to significant difficulty in analysis and losses in efficiency compared to the exact policy gradient. In sum, a model-free regret analysis remains a highly significant yet challenging open question due to the stochastic approximation type of algorithms involved.
@@
Recently, there has been some progress on regret analysis for continuous-time stochastic LQ RL in (Basei et al., , 2022; Szpruch et al., , 2024) that achieves sublinear regrets of their respective algorithms. Both papers assume that the diffusion coefficients are constant (independent of state and control), which is vital for their approaches to work. Again, in essence, these works are model-based because they apply either least-square or Bayesian methods to estimate model parameters and use the corresponding estimation errors to deduce the regret bounds. In particular, there is an intrinsic caveat of this approach applied to LQ problems: when estimating the model parameters, an unidentifiability issue may arise (from optimal control policies being linear feedbacks of the state) and a special exploratory stage has to be carefully designed to avoid it; see e.g., Szpruch et al., (2021); Basei et al., (2022); Szpruch et al., (2024); Xu et al., (2024). In addition, to get one estimate of the model parameters, the required number of episodes and that of timesteps within each episode increase exponentially over the learning process, adding substantial computational and memory costs. [^1] For example, given a total $N$ episodes of the learning budget, the algorithm in Basei et al., (2022) suggests to break it into several epoches, with the $k$-th epoch containing $2^{k}$ episodes and $2^{k/2}$ timesteps and to update a policy after having applied it for an epoch of episodes. As a result, although the total regret is proved to be $O(\log N)$, the storage and computational complexity increase exponentially in the number of epoches (which is the number of times of updating the policy parameters).
@@
In this paper, we propose a policy gradient based algorithm to solve a special class of continuous-time, finite-horizon stochastic LQ problems under the model-free, episodic RL setting, where the state processes are one dimensional and there is no running control award. Our main contributions are
@@
    We take a model-free approach to develop our algorithm, and base our analysis on the stochastic approximation scheme. In particular, the policy gradient in this paper is a “model-free gradient” instead of a “model-based gradient” commonly taken in discrete-time RL. As a result, we do not need to estimate model primitives in the entire analysis, circumventing the issues discussed earlier arising from estimating/learning those model parameters.
@@
The remainder of the paper is structured as follows. Section 2 formulates the problem and provides some preliminary results necessary for the subsequent development. Section 3 describes and explains the steps leading to our RL algorithm. Section 4 presents the main theoretical results on convergence and a regret bound of the proposed algorithm. Section 5 reports the results of numerical experiments. Finally, Section 6 concludes. The appendix contains the description of the numerical experiments and detailed proofs of statements are available at the full online version https://arxiv.org/pdf/2407.17226.
@@ 2.1 Classical Stochastic LQ Control
$\mathrm{d}x^{u}(t)=(Ax^{u}(t)+B^{\top}u(t))\mathrm{d}t+\sum_{j=1}^{m}(C_{j}x^{u}(t)+D_{j}^{\top}u(t))\mathrm{d}W^{(j)}(t),$ (2.1)
@@
$\max_{u}\mathbb{E}\left[\int_{0}^{T}-\frac{1}{2}Qx^{u}(t)^{2}\mathrm{d}t-\frac{1}{2}Hx^{u}(T)^{2}\right],$ (2.2)
@@
$V_{CL}(t,x)=\max_{u}\mathbb{E}\left[\int_{t}^{T}-\frac{1}{2}Qx^{u}(s)^{2}\mathrm{d}s-\frac{1}{2}Hx^{u}(T)^{2}\Big{|}x^{u}(t)=x\right].$ (2.3)
@@
$V_{CL}(t,x)=-\frac{1}{2}\biggl{[}\frac{Q}{\Lambda}+(H-\frac{Q}{\Lambda})e^{\Lambda(t-T)}\biggl{]}x^{2},$ (2.4)
$u_{CL}(t,x)=-(\sum_{j=1}^{m}D_{j}D_{j}^{\top})^{-1}(B+\sum_{j=1}^{m}C_{j}D_{j})x,$ (2.5)
@@
We remark that if $x$ is of a higher dimension and/or there is a control running reward, then the optimal policy will depend explicitly on the solution of the differential Riccati equation and hence become time-dependent (Yong and Zhou, , 1999, Chapter 6), for which our current method fails.  [^2] For example, when $x$ is higher dimensional, the optimal control is $u_{CL}(t,x)=-(\sum_{j=1}^{m}D_{j}^{\top}\frac{\partial V_{CL}(t,x)}{\partial xx}D_{j})^{-1}\left(B^{\top}\frac{\partial V_{CL}(t,x)}{\partial x}+\sum_{j=1}^{m}D_{j}^{\top}\frac{\partial V_{CL}(t,x)}{\partial xx}C_{j}x\right)$, which becomes time-dependent.
@@ 2.2 RL Theory for LQ
$\mathrm{d}x^{\pi}(t)=\widetilde{b}(x^{\pi}(t),\pi(\cdot,t))\mathrm{d}t+\sum_{j=1}^{m}\widetilde{\sigma}_{j}(x^{\pi}(t),\pi(\cdot,t))\mathrm{d}W^{(j)}(t),$ (2.6)
@@
$\widetilde{b}(x,\psi):=Ax+B^{\top}\int_{\mathbb{R}^{l}}u\psi(u)\mathrm{d}u,$ (2.7)
$\widetilde{\sigma}_{j}(x,\psi):=\sqrt{\int_{\mathbb{R}^{l}}(C_{j}x+D_{j}^{\top}u)^{2}\psi(u)\mathrm{d}u},\;\;(x,\psi)\in\mathbb{R}\times\mathcal{P}(\mathbb{R}^{l}).$ (2.8)
@@
$J(t,x;\pi)=\mathbb{E}\left[\int_{t}^{T}\left(-\frac{1}{2}Qx^{\pi}(s)^{2}+\gamma p^{\pi}(s)\right)\mathrm{d}s-\frac{1}{2}Hx^{\pi}(T)^{2}\Big{|}x^{\pi}(t)=x\right],$ (2.9)
@@
$V(t,x)=\max_{\pi}J(t,x;\pi).$ (2.10)
By standard stochastic control theory, the value function $V$ satisfies the HJB equation:
$V_{t}+\max_{\pi\in\mathcal{P}(\mathbb{R}^{l})}\biggl{\{}\int_{\mathbb{R}^{l}}\biggl{[}$ $V_{x}(Ax+Bu)+\frac{1}{2}\sum_{j=1}^{m}\left((C_{j}x+D_{j}u)^{2}V_{xx}\right)-\frac{1}{2}Qx^{2}$
$-\gamma\pi(u)\log\pi(u)\biggl{]}\pi(u)\mathrm{d}u\biggl{\}}=0,\quad V(T,x)=-\frac{1}{2}Hx^{2}.$
The verification theorem then yields the optimal policy
$\pi(u$ $\mid t,x)=$
$\mathcal{N}\left(u\Big{|}-(\sum_{j=1}^{m}D_{j}D_{j}^{\top}V_{xx})^{-1}(BV_{x}+\sum_{j=1}^{m}C_{j}D_{j}V_{xx}x),\gamma(\sum_{j=1}^{m}D_{j}D_{j}^{\top}V_{xx})^{-1}\right),$
where $\mathcal{N}(\cdot|\mu,\Sigma)$ is the multivariate Gaussian density with mean $\mu$ and covariance $\Sigma$. Plugging this back to the HJB equation, together with the ansatz
$V(t,x)=-\frac{1}{2}k_{1}(t)x^{2}+k_{3}(t)$ (2.11)
where $k_{1}>0$ and $k_{3}$ are certain functions of $t$, we obtain the optimal policy
$\pi(u\mid t,x)=\mathcal{N}\left(u\Big{|}-(\sum_{j=1}^{m}D_{j}D_{j}^{\top})^{-1}(B+\sum_{j=1}^{m}C_{j}D_{j})x,\frac{\gamma}{k_{1}(t)}(\sum_{j=1}^{m}D_{j}D_{j}^{\top})^{-1}\right).$ (2.12)
Moreover, the resulting HJB equation yields that $k_{1}$ and $k_{3}$ satisfy the following ODE:
$\frac{k_{1}^{\prime}(t)}{2}=-(A+B^{\top}\bar{\mu})k_{1}(t)-\frac{k_{1}(t)}{2}\sum_{j=1}^{m}\biggl{(}C_{j}^{2}+2C_{j}D_{j}^{\top}\bar{\mu}+D_{j}^{\top}\bar{\mu}\bar{\mu}^{\top}D_{j}\biggl{)}-\frac{Q}{2},\quad k_{1}(T)=H,$
and
$k_{3}^{\prime}(t)=\frac{k_{1}(t)}{2}\sum_{j=1}^{m}D_{j}^{\top}\Sigma D_{j}-\frac{\gamma}{2}\log\biggl{(}(2\pi e)^{l}\det(\Sigma)\biggl{)},\quad k_{3}(T)=0,$
where $\bar{\mu}=-(\sum_{j=1}^{m}D_{j}D_{j}^{\top})^{-1}(B+\sum_{j=1}^{m}C_{j}D_{j})$ and $\Sigma=\frac{\gamma}{k_{1}(t)}(\sum_{j=1}^{m}D_{j}D_{j}^{\top})^{-1}$.333 Detailed derivations of these results follow analogously from those in Wang et al., (2020) for the infinite horizon case. Note that these theoretical results cannot be used to compute the solution of the exploratory problem because the model parameters are unknown, yet they reveal the structure of the solution inherent to LQ RL (i.e. the optimal value function is quadratic in $x$ and optimal stochastic policy is Gaussian) that can be utilized to significantly reduce the complexity of function parameterization/approximation in learning.